OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = −1? (Please explain too) f(x) = −one fourth x^2 f(x) = one fourth x^2 f(x) = −4x^2 f(x) = 4x^2

hero (hero):

To find the equation of the parabola plug points (0,1) and (x, -1) into the formula \((x - x_1)^2 + (y - y_1)^2 = (x - x_2)^2 + (y - y_2)^2\) to get: \((x - 0)^2 + (y - 1)^2 = (x - x)^2 + (y - (-1))^2\) Then simplify: \(x^2 + (y - 1)^2 = 0 + (y + 1)^2\) And expand: \(x^2 + y^2 - 2y + 1 = y^2 + 2y + 1\) From here, continue isolating y.